OpenStudy (anonymous):

solve this equation, T/3 - 1/2 = 1 solve this equation, T/3 - 1/2 = 1 @Mathematics

OpenStudy (anonymous):

t/3=1+1/2=3/2 t=9/2

OpenStudy (anonymous):

8/9 - 1/5= a.31/45 b.7/45 c.1 22/44 d.1 3/4